# Formalization of Brownian motion in Lean

### v1

@@ 2 Formalization overview
This section gives an overview of the different theorems and definitions we formalized in order to construct Brownian motion. Subsequent sections will go into more detail for each of these steps. We refer to the implementation using links, like ProbabilityTheory.IsGaussian, which point either to Mathlib or to our repository. Our goal is to eventually integrate all of our code into Mathlib, and that process is ongoing. All links point to commit 82d756ea359cfa0f2c10bc00b8a997822c3b32ea for Mathlib, and commit aae12b7f25e401a52c60546a8ff0e3390e81b87b for our repository, which correspond to their state on November 22, 2025.
@@ 4.4.2 The main theorem
If furthermore $E$ is an extended metric space (not pseudo-metric: $d_{E}(x,y)=0$ implies $x=y$) and is complete, and $T$ is second-countable, then the process $X$ has a modification with Hölder continuous paths of exponent $\beta$ for all $\beta\in(0,(q-d)/p)$.
@@ 4.4.5 Building a Hölder continuous process
We discuss the proof of the second part of Theorem 4.20, which states that under additional assumptions on $E$ and $T$, the process $X$ has a modification with Hölder continuous paths. We first prove the result for a fixed $\beta\in(0,(q-d)/p)$, and will then deduce it for all $\beta$ in that interval simultaneously. An advantage of the formal proof is that it is easy to inspect the code to see which assumptions are used where. This part of the theorem requires $T$ to be second-countable because we first take a countable dense subset $T^{\prime}$ of $T$, on which we apply the main inequality of Theorem 4.20. We obtain that an expectation is finite, hence that the quantity inside the expectation is almost surely finite:
@@
The limits in that definition are the reason for the completeness assumption on $E$. Then $Y$ has Hölder continuous paths of order $\beta$. We finally need to show that $Y$ is a modification of $X$, which we do by showing that for all $t\in T$, $d_{E}(X_{t},Y_{t})=0$ almost surely. That is why that part of the theorem requires $E$ to be metric and not pseudo-metric. We refer to [KU23] for the proof that $Y$ is a modification of $X$, and only note that it uses convergence in probability, which we had to generalize in Mathlib from pseudo-metric to extended pseudo-metric spaces.
@@
Now that we have a modification $Y$ with Hölder continuous paths of order $\beta$ for a fixed $\beta\in(0,(q-d)/p)$, we want to obtain a modification with Hölder continuous paths of order $\gamma$ for all $\gamma\in(0,(q-d)/p)$ simultaneously. Since $T$ has finite diameter, being Hölder continuous of order $\beta$ implies being Hölder continuous of order $\gamma$ for all $\gamma\in(0,\beta]$ (which we proved at HolderOnWith.mono_right).
@@
Let $(\beta_{n})_{n\in\mathbb{N}}$ be an increasing sequence in (0,(qd)/p)(0,(q-d)/p) converging to (qd)/p(q-d)/p. For each $n$, we obtain a modification $Y^{n}$ with Hölder continuous paths of order $\beta_{n}$. Those $Y^{n}$ are all modifications of each other and are continuous, hence indistinguishable. Thus almost surely, for all $t\in T$, $Y^{n}_{t}=Y^{1}_{t}$ for all $n$. We can thus define a process $Z$ by $Z_{t}(\omega)=Y^{1}_{t}(\omega)$ if $\omega$ is in the event where all $Y^{n}$ are equal, and $Z_{t}(\omega)=x_{0}$ otherwise. Then $Z$ is a modification of $X$ and has Hölder continuous paths of order $\beta_{n}$ for all $n$. It thus has Hölder continuous paths of order $\gamma$ for all $\gamma\in(0,(q-d)/p)$.
@@
That discussion is very close to the one in [KU23], and we omitted an aspect that proved more subtle than expected during the formalization: the measurability of $Z$. The process $Z$ is defined as being equal to one of two measurable processes depending on whether $\omega$ belongs to a certain event or not. In order to ensure that $Z$ is measurable, that event must be measurable. In particular we need the event $\{\omega\mid\forall t\in T,\>Y_{t}^{1}(\omega)=Y_{t}^{n}(\omega)\}$ that two processes are equal to be measurable. We use continuity to reduce to equality of the processes on a countable subset of $T$, and it suffices then to prove that the set $\{\omega\mid Y_{t}^{1}(\omega)=Y_{t}^{n}(\omega)\}$ is measurable for all $t\in T$. Rewriting the equality as $d_{E}(Y_{t}^{1}(\omega),Y_{t}^{n}(\omega))=0$, we see that this is true if the distance $d_{E}(Y_{t}^{1},Y_{t}^{n})$ is a measurable function, which would follow from the measurability of the pairs $(Y_{t}^{1},Y_{t}^{n})$ with respect to $\mathcal{B}(E\times E)$. However, as currently stated, Theorem 4.20 does not guarantee that the two modifications $Y^{1}$ and $Y^{n}$ satisfy that property (unless $E$ is second-countable, in which case that follows from the measurability of each $Y^{i}$). The particular construction of $Y$ we used in the proof will be important to ensure the measurability: this phenomenon of relying on the particular object constructed in the proof rather than purely on the statement of the theorem that gives existence is not ideal but also not uncommon in mathematics and is an often encountered challenge in formalization [Mor25]. To solve it, we introduced the property for a process $Y$ of being a limit of $X$ on an event of probability one, and a fixed arbitrary value outside of that event (ProbabilityTheory.IsLimitOfIndicator), and our formal statement of Theorem 4.20 ensures that the modification we build satisfies that property. The fact that all $Y^{i}$ satisfy that property is sufficient to ensure the measurability of $Z$ and to prove that $Z$ itself has the same property, which can then be used in further constructions (needed for the localized Theorem 4.21).
@@ Acknowledgments
We would like to thank Markus Himmel for his many contributions to the formalization of the Kolmogorov-Chentsov proof. We are also grateful to Jonas Bayer, Lorenzo Luccioli, Alessio Rondelli, Jérémy Scanvic for their contributions to various parts of the formalization and to Pietro Monticone for his help with technical aspects of setting up the project.

### v2

@@ 2 Formalization overview
This section gives an overview of the different theorems and definitions we formalized in order to construct Brownian motion. Subsequent sections will go into more detail for each of these steps. We refer to the implementation using links, like ProbabilityTheory.IsGaussian, which point either to Mathlib or to our repository. Our goal is to eventually integrate all of our code into Mathlib, and that process is ongoing. All links point to commit 82d756ea359cfa0f2c10bc00b8a997822c3b32ea for Mathlib, and to commit 6a7a3141c6bc217484f635f3ab861316a37c0362 for our repository, which correspond its state on December 1, 2025.
@@ 4.4.2 The main theorem
If furthermore $E$ is complete and $T$ is second-countable, then the process $X$ has a modification with Hölder continuous paths of exponent $\beta$ for all $\beta\in(0,(q-d)/p)$.
@@ 4.4.5 Building a Hölder continuous process
We discuss the proof of the second part of Theorem 4.20, which states that for $E$ complete and $T$ second-countable, the process $X$ has a modification with Hölder continuous paths. We first prove the result for a fixed $\beta\in(0,(q-d)/p)$, and will then deduce it for all $\beta$ in that interval simultaneously. An advantage of the formal proof is that it is easy to inspect the code to see which assumptions are used where. This part of the theorem requires $T$ to be second-countable because we first take a countable dense subset $T^{\prime}$ of $T$, on which we apply the main inequality of Theorem 4.20. We obtain that an expectation is finite, hence that the quantity inside the expectation is almost surely finite:
@@
The limits in that definition are the reason for the completeness assumption on $E$. Then $Y$ has Hölder continuous paths of order $\beta$ and is measurable. We finally need to show that $Y$ is a modification of $X$. It is actually only a “distance zero modification”, by which we mean that $d_{E}(X_{t},Y_{t})=0$ almost surely. Since $E$ is only a pseudo-metric space, $Y$ might not satisfy the modification property $Y_{t}=X_{t}$ almost surely. We refer to [KU23] for the proof that $Y$ is a distance zero modification of XX [^6] They use a metric space assumption for $E$ and prove that $Y$ is a true modification: replacing equalities with distance zero statements gives our proof for pseudo-metric spaces., and only note that it uses convergence in probability, which we had to generalize in Mathlib from pseudo-metric to extended pseudo-metric spaces. From a distance zero modification $Y$ of $X$, we can obtain a true modification in a pseudo-metric space as follows: for each $t\in T$, let $Z_{t}$ be equal to $X_{t}$ on the event where $d_{E}(X_{t},Y_{t})=0$, and equal to $Y_{0}$ otherwise. Then $Z_{t}=X_{t}$ almost surely, and since $d_{E}(Z_{t}(\omega),Y_{t}(\omega))=0$ for all $t$ and all $\omega$, $Z$ has the same continuity properties as $Y$. Finally, $Z$ is measurable (that is, $Z_{t}$ is measurable for all $t$), since a random variable that is at distance zero from a measurable random variable is itself measurable.
@@
Now that we have a modification with Hölder continuous paths of order $\beta$ for a fixed $\beta\in(0,(q-d)/p)$, we want to obtain a modification with Hölder continuous paths of order $\gamma$ for all $\gamma\in(0,(q-d)/p)$ simultaneously. Since $T$ has finite diameter, being Hölder continuous of order $\beta$ implies being Hölder continuous of order $\gamma$ for all $\gamma\in(0,\beta]$ (which we proved at HolderOnWith.mono_right).
@@
Let $(\beta_{n})_{n\in\mathbb{N}}$ be an increasing sequence in (0,(qd)/p)(0,(q-d)/p) converging to (qd)/p(q-d)/p. For each $n$, we obtain a modification $Y^{n}$ with Hölder continuous paths of order $\beta_{n}$. Those $Y^{n}$ are all distance zero modifications of each other and are continuous, hence indistinguishable (in the sense of having distance zero). Thus almost surely, for all $t\in T$, $d_{E}(Y^{n}_{t},Y^{1}_{t})=0$ for all $n$. We can thus define a process $Z$ by $Z_{t}(\omega)=Y^{1}_{t}(\omega)$ if $\omega$ is in the event where all $Y^{n}$ are at distance zero, and $Z_{t}(\omega)=x_{0}$ otherwise. Then $Z$ is a modification of $X$ and has Hölder continuous paths of order $\beta_{n}$ for all $n$. It thus has Hölder continuous paths of order $\gamma$ for all $\gamma\in(0,(q-d)/p)$.
@@
That discussion is very close to the one in [KU23] (up to our use of distance zero instead of equality), and we omitted an aspect that proved more subtle than expected during the formalization: the measurability of $Z$. The process $Z$ is defined as being equal to one of two measurable processes depending on whether $\omega$ belongs to a certain event or not. In order to ensure that $Z$ is measurable, that event must be measurable. In particular we need the event $\{\omega\mid\forall t\in T,\>d_{E}(Y_{t}^{1}(\omega),Y_{t}^{n}(\omega))=0\}$ that two processes are at distance zero to be measurable. We use continuity to reduce to a countable subset of $T$, and it suffices then to prove that the set $\{\omega\mid d_{E}(Y_{t}^{1}(\omega),Y_{t}^{n}(\omega))=0\}$ is measurable for all $t\in T$. This is true if the distance $d_{E}(Y_{t}^{1},Y_{t}^{n})$ is a measurable function, which would follow from the measurability of the pairs $(Y_{t}^{1},Y_{t}^{n})$ with respect to $\mathcal{B}(E\times E)$. However, as currently stated, Theorem 4.20 does not guarantee that the two modifications $Y^{1}$ and $Y^{n}$ satisfy that property (unless $E$ is second-countable, in which case that follows from the measurability of each $Y^{i}$). The particular construction of $Y$ we used in the proof will be important to ensure the measurability: this phenomenon of relying on the particular object constructed in the proof rather than purely on the statement of the theorem that gives existence is not ideal but also not uncommon in mathematics and is an often encountered challenge in formalization [Mor25]. To solve it, we introduced the property for a process $Y$ of being at distance zero of a limit of $X$ on an event of probability one, and at distance zero of a fixed arbitrary value outside of that event (ProbabilityTheory.IsLimitOfIndicator), and our formal statement of Theorem 4.20 ensures that the modification we build satisfies that property. The fact that all $Y^{i}$ satisfy that property is sufficient to ensure the measurability of $Z$ and to prove that $Z$ itself has the same property, which can then be used in further constructions (needed for the localized Theorem 4.21).
@@ Acknowledgments
We would like to thank Markus Himmel for his many contributions to the formalization of the Kolmogorov-Chentsov proof, as well as Sébastien Gouëzel, whose suggestions helped us greatly improve Theorem 4.20 by removing any need for a metric space (instead of pseudo-metric space) assumption. We are also grateful to Jonas Bayer, Lorenzo Luccioli, Alessio Rondelli, Jérémy Scanvic for their contributions to various parts of the formalization and to Pietro Monticone for his help with technical aspects of setting up the project.